# Soft-TransFormers for Continual Learning

### v2

@@ Abstract
Inspired by Well-initialized Lottery Ticket Hypothesis (WLTH), which provides suboptimal fine-tuning solutions, we propose a novel fully fine-tuned continual learning (CL) method referred to as Soft-TransFormers (Soft-TF). Soft-TF sequentially learns and selects an optimal soft-network for each task. During sequential training in CL, an well-initialized Soft-TF mask optimizes the weights of sparse layers to obtain task-adaptive soft (real-valued) networks, while keeping the well-pre-trained layer parameters frozen. In inference, the identified task-adaptive network of Soft-TF masks the parameters of the pre-trained network, mapping to an optimal solution for each task and minimizing Catastrophic Forgetting (CF) - the soft-masking preserves the knowledge of the pre-trained network. Extensive experiments on the Vision Transformer (ViT) and the Language Transformer (Bert) demonstrate the effectiveness of Soft-TF, achieving state-of-the-art performance across Vision and Language Class Incremental Learning (CIL) scenarios.
Keywords: Continual Learning (CL), Soft-TransFormers (Soft-TF), Adapter, and LoRA.
@@ 1 Introduction
Continual Learning (CL), also known as Lifelong Learning (Thrun, 1995; Rusu et al., 2016; Zenke et al., 2017; Hassabis et al., 2017), is a learning paradigm where a series of tasks are learned sequentially. The principle objective of continual learning is to replicate human cognition, characterized by the ability to learn new concepts incrementally throughout one’s lifespan. An optimal continual learning system could facilitate a positive forward and backward transfer, leveraging the knowledge gained from previous tasks to solve new ones, while also updating its understanding of previous tasks with the new knowledge. However, achieving continual learning is challenging due to the occurrence of catastrophic forgetting or catastrophic interference (McCloskey and Cohen, 1989), a phenomenon where the performance of the model on previous tasks deteriorates significantly when it learns new tasks. This can make it challenging to retain the knowledge acquired from previous tasks, ultimately leading to a decrease in overall performance. To address the issue of catastrophic forgetting during continual learning, numerous conventional approaches have been proposed on Convolutional Neural Networks (CNNs), which can be broadly classified as follows: (1) Regularization-based methods (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020; Titsias et al., 2020; Mirzadeh et al., 2021) aim to keep the learned information of past tasks during continual training aided by sophisticatedly designed regularization terms, (2) Rehearsal-based methods (Rebuffi et al., 2017; Riemer et al., 2018; Chaudhry et al., 2019a, b; Saha et al., 2021) utilize a set of real or synthesized data from the previous tasks and revisit them, and (3) Architecture-based methods (Mallya et al., 2018; Serrà et al., 2018; Li et al., 2019; Wortsman et al., 2020; Kang et al., 2022, 2023, 2024a) propose to minimize the inter-task interference via newly designed architectural components.
@@
Developing neural network models that leverage large-scaled pre-trained models. i.e., Vision Transformer (ViT) (Dosovitskiy et al., 2020) and Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) leads to a new paradigm shift referred to as (4) Prompt-based methods in Continual Learning (CL). Prompt-based methods learn continual representations to provide fixed pre-trained transformers with additional instruction. Notably, while L2P (Wang et al., 2022c) stands out as the seminal work that bridges the gap between prompting and continual learning, DualPrompt (Wang et al., 2022b) introduces an innovative approach to affixing complementary prompts to the pre-trained backbone, thereby enabling the acquisition of both task-invariant and task-specific instructions. Additionally, other notable contributions in this field encompass DyTox (Douillard et al., 2022), S-Prompt (Wang et al., 2022a), CODA-P (Smith et al., 2023b), ConStruct-VL (Smith et al., 2023a), ST-Prompt (Pei et al., 2023), and LGCL (Khan et al., 2023). Recently, Qiao et al. (2024) investigated prompt-projection for better generalized continual learners.
With prior developments of representational research, current prompt-based models can be fine-tuned using trainable prompts to improve their performance on sequential tasks, and the fixed pre-trained backbone can consistently provide unforgettable base session knowledge. However, prompt-based models come with several disadvantages and limitations. First, the prompt-based CL model is highly dependent on the quality and design of the sample or the task-relevant prompts. Poorly trained prompts could lead to suboptimal performance or tend to be biased. Second, managing and maintaining a large set of prompts can become cumbersome and unmanageable as the number of tasks increases. Lastly, prompt tuning is not as flexible as full fine-tuning. The only prompt-tuning with the pre-trained model cannot capture all the nuances of uncorrelated sequential tasks even though leveraging the frozen well-initialized model pre-trained on large-scale datasets since the only frozen well-initialized model provides global solutions rather than task-specific solutions. These disadvantages motivate us to make informed decisions about when and how to use prompt-based models and explore alternative methods like full fine-tuning for more robust and flexible prompt-based continual learning performance.
To overcome the limitations of conventional prompt-based methods, the central focus of this work is to pinpoint the parameter-efficient fine-tuning representations of frozen pre-trained networks, comparing with conventional LLM-based fine-tuning methods such as Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2021), as shown in Figure 2. We focus on two main issues when sequential full fine-tuning the pre-trained foundation models: (1) Catastrophic Forgetting (CF) and (2) parameter-efficient fine-tuning CL model. To deploy a practical model to deal with the two points, we suggest a new paradigm for Continual Learning (CL), named Well-initialized Lottery Ticket Hypotehesis:
Well-initialized Lottery Ticket Hypothesis (WLTH). A well-initialized dense neural network contains globally minimal solutions that can retain the prior class knowledge while providing room to learn the new class knowledge through isolated fine-tuning of the networks or subnetworks.
Leveraged by the WLTH, this work proposes a new Soft-TransFormer (Soft-TF) to address fine-tuning with minimal CF, as shown in Figure 1. In this work, we find task-specific soft-networks based on well-trained frozen transformer parameters that incrementally learn task-adaptive weights associated with sequential tasks.
Our contributions can be summarized as follows:
@@
  • Inspired by Well-initialized Lottery Ticket Hypothesis (WLTH), we propose a novel continual learning method referred to as Soft-TransFormers (Soft-TF), which learns compact task-specific soft-networks from well pre-trained parameters for each task.
  • Extensive experiments demonstrate that Soft-TF leads to better generalized parameter-efficient continual models than baselines such as DualPrompts, Adaptor, and LoRA, achieving state-of-the-art performances on Vision and Language Class-Incremental Learning (CIL) scenarios.
@@
(a) Adapter (b) LoRA (c) Soft-TransFormer
@@
Figure 2: Comparisions of LLM-based fine-tuning methods (Adapters, LoRA) with Soft-TransFormer (Soft-TF): well-initialized Soft-TF $\bm{\theta}\odot\bm{m}_{t}$ is trained at $l_{e}$-th or only few attention layers while LLM fine-tuning methods such as Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2021) are trained at all attention layers. Note that all pre-trained parameters $\bm{\theta}$ are fixed in fine-tuning t-th task parameter $\bm{m}_{t}$ in Continual Learning (CL) scenarios.
2 Related Works
Continual Learning (McCloskey and Cohen, 1989; Thrun, 1995; Kumar and Daume III, 2012; Li and Hoiem, 2016) is the challenge of learning a sequence of tasks continuously while utilizing and preserving previously learned knowledge to improve performance on new tasks. Four major approaches have been proposed to tackle the challenges of continual learning, such as catastrophic forgetting. One such approach is (1) Regularization-based approaches (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020; Titsias et al., 2020; Mirzadeh et al., 2021), which aim to reduce catastrophic forgetting by imposing regularization constraints that inhibit changes to the weights or nodes associated with past tasks. (2) Rehearsal-based approaches (Rebuffi et al., 2017; Chaudhry et al., 2019a, b; Saha et al., 2021; Deng et al., 2021; Sun et al., 2023; Sarfraz et al., 2023; Mai et al., 2021; Lin et al., 2023; Aljundi et al., 2019; Caccia et al., 2021; Chaudhry et al., 2019c; Liang and Li, 2024; Buzzega et al., 2020) store small data summaries to the past tasks and replay them during training to retain the acquired knowledge. Some approaches in this line of work (Shin et al., 2017; Aljundi et al., 2019) accommodate the generative model to construct the pseudo-rehearsals for previous tasks. (3) Architecture-based approaches (Mallya et al., 2018; Serrà et al., 2018; Li et al., 2019; Wortsman et al., 2020; Kang et al., 2022, 2023, 2024b, 2024a) use the additional capacity to expand (Xu and Zhu, 2018), dynamic representation (Yan et al., 2021; Singh et al., 2020) or isolate (Rusu et al., 2016) model parameters, preserving learned knowledge and preventing forgetting. Rehearsal and architecture-based methods have shown remarkable efficacy in suppressing catastrophic forgetting but require additional capacity for the task-adaptive parameters (Wortsman et al., 2020) or the replay buffers. Recently, (4) Prompt-based approaches, an emerging transfer learning technique, harnesses a fixed function of pre-trained Transformer models. This empowers the language model to receive additional instructions for enhancing its performance on downstream tasks. Two key works stand out in the prompt-based CL: L2P (Wang et al., 2022c), a seminal study bridging the gap between prompting and continual learning, and DualPrompt (Wang et al., 2022b), which proposes an innovative method for attaching complementary prompts to a fixed pre-trained backbone. Here, for the first time, we introduce a new approach to update the fixed pre-trained parameters through learnable sparse layered soft-networks under the convergence theory, maximumly enabling the acquisition of task-invariant and task-specific instructions with prompts.
Prompt-based CL. With advances in Vision Transformers (Khan et al., 2022) and prompt-based fine-tuning in NLP (Li and Liang, 2021) and CL (Razdaibiedina et al., 2023), Wang et al. (2022c) have shown that interacting with a pre-trained model via prompt learning is a promising approach, L2P  (Wang et al., 2022c) for continual learning. Then, other advanced prompt-based CL models emerged such as DualPrompt (Wang et al., 2022b), DyTox (Douillard et al., 2022), S-Prompt (Wang et al., 2022a), CODA-P (Smith et al., 2023b), ConStruct-VL (Smith et al., 2023a), ST-Prompt (Pei et al., 2023), and LGCL (Khan et al., 2023). Recently, Prompt Gradient Projection (PGP) (Qiao et al., 2024), a small set of learnable orthogonal parameters, is appended to the input and enables quick adaptation of a frozen ImageNet pre-trained model to new streaming tasks. Their analysis shows that directly leveraging the pre-trained vision-language model without introducing any learnable parameters is a simple yet promising approach to continual learning. The PGP was also adopted to a joint vision-language model like CLIP (Radford et al., 2021) for continual learning, which presents multiple advantages - catering to practical scenarios without well-defined task identities and boundaries in class incremental learning scenarios. However, prompt-based models come with several disadvantages. Poorly trained prompts could lead to suboptimal performance or tend to be biased. Moreover, prompt tuning could not capture all nuances of uncorrelated sequential tasks. These disadvantages lead to exploring alternative methods like full fine-tuning or hybrid approaches for more robust and flexible prompt-based model performance. In this work, to alleviate these issues of prompt tuning, we investigate a fully fine-tuning of well-pre-trained transformers on training soft networks on various continual learning scenarios, comparing with LLM-based CL baselines such as MBPA++ (de Masson D’Autume et al., 2019), IDBR (Huang et al., 2021), and SLM (Bohao et al., 2024).
@@ 3 Prerequisites
We start with conventional prompt-based continual learning methods using Vision Transformer (ViT) (Dosovitskiy et al., 2020) in Class Incremental Learning (CIL) scenarios.
@@ 3.2 Prompt-based Class Incremental Learning (CIL)
A simple yet effective prompt-based (prompt-tuning) CIL model: Learning to Prompt (L2P) (Wang et al., 2022c) is first proposed. In this model, a prompt $p$, a tiny set of trainable tokens combined with image features, is fed into the Vision Transformer (ViT) to help the model resist forgetting. To select suitable prompts for task-specific training, L2P utilizes a prompt pool $P$ containing numerous prompt-key pairs, $\{p_{t},k_{t}\}_{t=1}^{\mathcal{T}}$, where ${p}_{t}\in\mathbb{R}^{1\times D}$ represents the $t$-th task prompt, $k_{t}$ represents the $t$-th coresponding task key, and $\mathcal{T}$ is the total number of prompt-key pairs. Building on L2P, DualPrompt (Wang et al., 2022b) divided the prompts into expert (E-) prompts and general (G-) prompts for distinct features learning. DualPrompt also replaced prompt-tuning with prefix-tuning, which was successfully proven in Language CIL scenarios.
4 Transformer with Learnable Soft-networks
In this section, we explain how Soft-TransFormers (Soft-TF) leverage learnable soft-networks to train sequential tasks while keeping the well-pretrained model parameters fixed. To introduce our novel Soft-TF and provide a clearer understanding, we draw on a partial explanation of DualPrompt.

### v3

# Soft-TransFormer for Continual Learning

@@ Abstract
Inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH), we introduce Soft-Transformer (Soft-TF), a parameter-efficient framework for continual learning that leverages soft, real-valued subnetworks over a frozen pre-trained Transformer. Instead of relying on manually designed prompts or adapters, Soft-TF learns task-specific multiplicative masks applied to the key, query, value, and output projections in self-attention. These masks enable smooth and stable task adaptation while preserving shared representations. Combined with a lightweight dual-prompt mechanism, Soft-TF maintains strong knowledge retention and mitigates Catastrophic Forgetting (CF). Across multiple continual learning benchmarks, Soft-TF achieves state-of-the-art performance, consistently outperforming prompt-based, adapter-based, and LoRA-style baselines while requiring minimal additional parameters.
Machine Learning, Continual Learning, Sparse TransFormers
@@ 1 Introduction
Continual Learning (CL), also known as Lifelong Learning (Thrun, 1994; Rusu et al., 2016; Zenke et al., 2017; Hassabis et al., 2017), aims to endow models with the ability to acquire new knowledge over time without erasing previously learned information. An ideal CL system would enable both forward transfer, where prior knowledge facilitates learning new tasks, and backward transfer, where new experiences refine earlier representations. However, modern neural networks remain highly susceptible to catastrophic forgetting (McCloskey and Cohen, 1989), where updates from new tasks interfere with parameters essential for earlier ones, resulting in significant performance degradation.
A large body of prior work - predominantly developed for convolutional neural networks - tackles catastrophic forgetting through three major strategies: (1) Regularization-based methods (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020; Titsias et al., 2020; Mirzadeh et al., 2021) constrain updates to preserve knowledge of previously learned tasks; (2) Rehearsal-based methods (Rebuffi et al., 2017; Riemer et al., 2018; Chaudhry et al., 2019a, b; Saha et al., 2021) store or synthesize past data for replay; and (3) Architecture-based methods (Mallya et al., 2018; Serrà et al., 2018; Li et al., 2019; Wortsman et al., 2020; Kang et al., 2022) reduce inter-task interference via task-specific modules or structural expansions.
@@
With the advent of large-scale pre-trained Transformers, i.e., ViT (Dosovitskiy et al., 2021), a new paradigm - (4) Prompt-based Continual Learning—has emerged. By freezing the backbone and learning lightweight prompts, these methods adapt the model to sequential tasks while preserving the strong representational priors of the pre-trained architecture. L2P (Wang et al., 2022d) first demonstrated that prompt retrieval can bridge prompting and continual representation learning. DualPrompt (Wang et al., 2022c) further decomposes prompts into task-general and task-specific components, improving adaptation and stability. Subsequent advances include DyTox (Douillard et al., 2022), S-Prompt (Wang et al., 2022a), CODA-P (Smith et al., 2023b), ConStruct-VL (Smith et al., 2023a), ST-Prompt (Pei et al., 2023), LGCL (Khan et al., 2023), ESN (Wang et al., 2022b), CPrompt (Gao et al., 2024), and NoRGa (Le et al., 2024).
Despite their effectiveness, prompt-based CL methods exhibit several important limitations. First, performance depends heavily on the quality, granularity, and design of prompts, making the method sensitive to prompt initialization and capacity. Second, as the number of tasks grows, managing a large and diverse prompt pool becomes cumbersome. Third, prompt tuning inherently lacks the expressiveness of full fine-tuning-restricted prompts alone cannot capture the fine-grained task-specific variations needed for diverse and uncorrelated sequential tasks. While frozen pre-trained models offer strong global priors, they do not directly provide task-adaptive pathways, limiting the flexibility of prompt-only tuning. These limitations motivate the exploration of alternative fine-tuning paradigms that combine the stability of frozen backbones with the flexibility of full parameter adaptation.
To address these challenges, we revisit the parameter-efficient fine-tuning landscape, especially in comparison to methods such as Adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2022), as illustrated in Figure˜2. We focus on two key desiderata in sequentially adapting pre-trained foundation models: (1) minimizing catastrophic forgetting and (2) ensuring parameter-efficiency during continual adaptation. Motivated by these goals, we introduce the following hypothesis:
Well-initialized Lottery Ticket Hypothesis (WLTH). A well-initialized dense neural network contains structured, globally stable solutions that preserve prior task knowledge, while also exposing local subspaces that can be selectively adapted for new tasks. These subspaces - corresponding to soft, real-valued subnetworks - enable isolated and parameter-efficient fine-tuning without overwriting the information encoded in the frozen backbone.
Inspired by WLTH, we propose Soft-Transformer (Soft-TF), a new parameter-efficient continual learning framework that discovers task-specific soft subnetworks within well-trained Transformer backbones, as shown in Figure˜1. Soft-TF incrementally learns task-adaptive multiplicative masks that modulate key, query, value, and output projections in self-attention layers, enabling flexible and stable adaptation with minimal interference between tasks.
Our main contributions are as follows:
@@
  • We propose Soft-Transformer (Soft-TF), a new continual learning paradigm inspired by the Well-initialized Lottery Ticket Hypothesis, which learns compact task-specific soft subnetworks from well pre-trained Transformers.
  • We demonstrate that Soft-TF achieves state-of-the-art parameter-efficient continual learning performance, outperforming DualPrompt, NoRGa, Adapter, and LoRA across class-incremental learning scenarios.
@@
(a) Adapter (b) LoRA (c) our Soft-TransFormer
@@
Figure 2: Comparisions of LLM-based fine-tuning methods (Adapters, LoRA) with Soft-TransFormer (Soft-TF): well-initialized Soft-TF $\bm{\theta}\odot\bm{m}_{t}$ is trained at $l_{e}$-th or only few attention layers while LLM fine-tuning methods such as Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2022) are trained at all attention layers. Note that all pre-trained parameters $\bm{\theta}$ are fixed in fine-tuning t-th task parameter $\bm{m}_{t}$ in Continual Learning (CL) scenarios.
2 Related Work
Continual Learning. Continual Learning (CL) aims to enable models to accumulate knowledge over a sequence of tasks while preserving performance on previously learned ones (McCloskey and Cohen, 1989; Thrun, 1994). Early efforts focused on mitigating catastrophic forgetting, the tendency of neural networks to override earlier representations when trained on new data. Regularization-based methods address this by constraining parameter updates that would negatively impact past tasks, often by estimating parameter importance or imposing functional stability (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020). Rehearsal-based approaches alleviate forgetting through explicit memory replay, either by storing exemplar sets or synthesizing pseudo-data from generative models (Rebuffi et al., 2017; Chaudhry et al., 2019a; Buzzega et al., 2020). Architecture-based strategies expand or isolate parts of the model so that new tasks use disjoint or sparsely overlapping parameter subsets, thereby preventing interference (Mallya et al., 2018; Serrà et al., 2018; Wortsman et al., 2020; Kang et al., 2022). Although these methods have proven effective, they often require additional storage, dynamic model growth, or explicit replay buffers, which can become impractical in long-horizon or resource-constrained CL scenarios.
Parameter-efficient Adaptation. The shift from training models from scratch toward adapting large-scale pre-trained Transformers - such as ViT and CLIP (Dosovitskiy et al., 2021; Radford et al., 2021) - has significantly influenced recent CL research. Parameter-efficient fine-tuning (PEFT) methods, including adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2022), introduce compact trainable modules while keeping most backbone parameters frozen. These approaches provide stronger adaptation capacity than traditional CL techniques but still accumulate per-task modules or suffer from interference across tasks when task similarity is low. Prompt-based CL offers an even lighter adaptation mechanism: instead of modifying model weights, prompts act as learnable soft instructions prepended or inserted into the Transformer architecture. L2P (Wang et al., 2022d) initiated this direction by retrieving task-relevant prompts from a pool, while DualPrompt (Wang et al., 2022c) decomposed prompts into task-general and task-specific components to balance stability and plasticity. More recent advances aim to improve the expressiveness and robustness of prompt-based continual learning. CODA-Prompt (Smith et al., 2023b) utilizes decomposed attention to assemble input-conditioned prompts, mitigating task interference, while S-Prompt (Wang et al., 2022a) emphasizes simplicity and stability via domain-independent prompts. HiDe-Prompt (Wang et al., 2023) explores hierarchical and disentangled prompt representations to better capture task structures, and NoRGa (Le et al., 2024) introduces non-linear residual gating to enhance adaptability by connecting prompting with mixture-of-experts mechanisms. Collectively, these methods highlight the potential of prompt-based PEFT for continual learning, while also exposing inherent limitations in relying solely on prompt-level conditioning for scalable and expressive task adaptation.
Limitations of Prompt-based CL. Despite their attractive efficiency, prompt-based methods face intrinsic constraints that limit their robustness in practical continual learning settings. Since prompts modify representations only implicitly through input conditioning, their expressiveness is fundamentally lower than module-level or weight-level adaptation, making them less suited for highly heterogeneous or weakly correlated task sequences. Moreover, prompt retrieval and initialization are sensitive to distribution shifts, and prompt pools may grow without bound as tasks accumulate, reducing scalability and increasing memory cost. Adapter- and LoRA-based PEFT methods provide stronger adaptation capacity but may introduce task interference or undesirable parameter growth. These challenges highlight the need for a principled and expressive mechanism that maintains the efficiency of PEFT while allowing direct modulation of the pre-trained backbone. Motivated by the Well-initialized Lottery Ticket Hypothesis (WLTH), we propose a Soft-Transformer (Soft-TF), which learns soft, real-valued subnetworks that multiplicatively modulate attention projections in a frozen Transformer. By selectively activating task-relevant subspaces within a well-initialized model, Soft-TF achieves stable, expressive, and parameter-efficient continual adaptation without relying solely on prompt proliferation.
Figure 3: Soft-TransFormer (Soft-TF): At training time, the E-Prompt and the Soft-network are selected according to task identity, and the selected G-Prompt, E-Prompt, and the Soft-MSAs are trained together with a classifier. At test time, an input is transformed by a query function (Prompt ID) or task identifier to match the closest task key $\bm{k}_{t}$, E-prompt $\bm{e}_{t}$ and Soft-networks $\bm{m}_{t}^{\{K,Q,V\}}$.
@@ 3 Prerequisites
We start with conventional prompt-based continual learning methods using Vision Transformer (ViT) (Dosovitskiy et al., 2021) in Class Incremental Learning (CIL) scenarios.
@@ 3.2 Prompt-based Class Incremental Learning (CIL)
A simple yet effective prompt-based (prompt-tuning) CIL model: Learning to Prompt (L2P) (Wang et al., 2022d) is first proposed. In this model, a prompt $\bm{p}$, a tiny set of trainable tokens combined with image features, is fed into the Vision Transformer (ViT) to help the model resist forgetting. To select suitable prompts for task-specific training, L2P utilizes a prompt pool $\bm{P}$ containing numerous prompt-key pairs, $\{\bm{p}_{t},\bm{k}_{t}\}_{t=1}^{\mathcal{T}}$, where $\bm{p}_{t}\in\mathbb{R}^{1\times D}$ represents the $t$-th task prompt, $\bm{k}_{t}$ represents the $t$-th coresponding task key, and $\mathcal{T}$ is the total number of prompt-key pairs. Building on L2P, DualPrompt (Wang et al., 2022c) divided the prompts into expert (E-) prompts and general (G-) prompts for distinct features learning. DualPrompt also replaced prompt-tuning with prefix-tuning, which was successfully proven in CIL scenarios.
4 Soft-TransFormer
This section presents Soft-TransFormer (Soft-TF), a parameter-efficient continual learning framework that adapts a frozen pre-trained Transformer through continuous, learnable soft-subnetworks and structured multi-layer prompting. Soft-TF preserves the rich representational structure encoded in the pre-trained backbone while enabling flexible task-specific modulation with minimal computational overhead and significantly reduced forgetting.